# Fisher-Geometric Diffusion in Stochastic Gradient Descent: Optimal Rates, Oracle Complexity, and Information-Theoretic Limits

### 10 Apr 2026 v3

@@ 3.7 Key geometric parameters (preview)
4 Fisher/Godambe Alignment of Mini-Batch Noise
@@ 4.1 Finite-population covariance identity (exact, conditional on data)
4.2 From exchangeability to Godambe/Fisher alignment (population geometry)
@@
Theorem 4.3 (Godambe alignment under exchangeability).
@@
Interpretation. Equation (22) is exact (not an approximation) when one marginalizes over both the mini-batch index $B$ and the dataset realization conditional on $\mu$. The finite-population correction (1b/n)(1-b/n) of Lemma 4.1 is present when one conditions on a fixed dataset, but it disappears at the population level because the randomness of the dataset itself contributes a complementary term through the total-covariance decomposition. Equation (23) below makes this bookkeeping explicit.
@@
Figure 1: Isotropic noise (left) vs. Fisher-aligned noise (right) at the same total variance. Batch size rescales temperature but not shape.
@@
5 Continuous-Time Diffusion Approximation and the OU Regime
This section links the discrete-time SGD recursion to a continuous-time SDE whose diffusion covariance is determined by the mini-batch sampling model and the intrinsic noise geometry derived in Section 4. Specifically, $G^{\star}(\theta)$ was defined in Section 3 and shown in Theorem 4.3 to govern the leading-order mini-batch covariance. Near a nondegenerate critical point, the SDE linearizes to an Ornstein–Uhlenbeck (OU) process, and its stationary covariance is characterized by a Lyapunov equation. The resulting “temperature” parameter $\tau=\eta/b$ is the scalar that converts sampling effort into diffusion strength. Diffusion approximations are a standard tool in OR for reducing complex stochastic systems to tractable Brownian models (Harrison, 1985; Whitt, 2002); the present setting is analogous, with the sampling mechanism playing the role of the arrival process.
The stationary covariance and risk expressions derived below depend only on the local second-moment structure of the oracle noise and do not require independence across oracle calls; a martingale-difference oracle (Assumption 6.9) suffices.
5.1 SGD decomposition and effective temperature
Fix constant stepsize $\eta>0$ and batch size $b$. Let
@@ 5.1 SGD decomposition and effective temperature
and write the mini-batch gradient estimator as
$g_{t}(\theta_{t})=h(\theta_{t})+\xi_{t}(\theta_{t}),\qquad\mathbb{E}[\xi_{t}(\theta_{t})\mid\mathcal{F}_{t}]=0,$
where $\mathcal{F}_{t}$ is the natural filtration. Then SGD is
$\theta_{t+1}=\theta_{t}-\eta\,h(\theta_{t})-\eta\,\xi_{t}(\theta_{t}).$ (24)
By Theorem 4.3 (and its Fisher specialization in Remark 4.7), the conditional covariance of the noise obeys
$\mathrm{Cov}\!\bigl(\xi_{t}(\theta)\mid\mathcal{F}_{t}\bigr)=\frac{1}{b}\,G^{\star}(\theta)+r_{b,n}(\theta),\qquad\|r_{b,n}(\theta)\|_{\operatorname{op}}=o(b^{-1})+O(n^{-1}),$ (25)
uniformly in a local neighborhood (the $O(n^{-1})$ term accounts for finite-population corrections and covariance-estimation bias; it is negligible whenever $b=o(n)$).
We call $\tau:=\eta/b$ the effective temperature:
@@
which will appear as the single scalar factor multiplying the diffusion covariance in the continuous-time limit.
5.2 A diffusion approximation (Euler–Maruyama viewpoint)
Let $s=t\eta$ denote algorithmic time and define the piecewise-linear interpolation $\bar{\theta}^{\eta}(\cdot)$ by $\bar{\theta}^{\eta}(t\eta)=\theta_{t}$ and linearity on $t\eta,(t+1)\eta]$.
Assumption 5.1 (Local regularity and localization).
On a neighborhood $\mathcal{U}$ of $\theta^{\star}$:
@@ 5.2 A diffusion approximation (Euler–Maruyama viewpoint)
    $h(\theta):=\nabla L(\theta)$ is locally Lipschitz;
@@
    $\nabla^{2}L(\theta)$ is locally Lipschitz (i.e., there exists $L_{H}<\infty$ such that $\|\nabla^{2}L(\theta)-\nabla^{2}L(\theta^{\prime})\|_{\operatorname{op}}\leq L_{H}\|\theta-\theta^{\prime}\|_{2}$ for all $\theta,\theta^{\prime}\in\mathcal{U}$);
@@
    $\sup_{\theta\in\mathcal{U}}\mathbb{E}\|\psi(\theta;Z)\|^{4}<\infty$ (bounded fourth moment; implied by Assumption 3.8(b));
  4. (iv)
    $H^{\star}:=\nabla^{2}L(\theta^{\star})\succ 0$ (local strong convexity; Assumption 3.9);
  5. (v)
    for each fixed horizon $T>0$, $\Pr(\theta_{t}\in\mathcal{U}\ \forall\,t\leq T/\eta)\to 1$ as $\eta\downarrow 0$ (localization; standard in stochastic approximation).
Assumption 5.2 (Asymptotic conditional covariance and Lindeberg).
Uniformly on $\mathcal{U}$, $\mathrm{Cov}(\xi_{t}(\theta)\mid\mathcal{F}_{t})=b^{-1}G^{\star}(\theta)+o(b^{-1})$ and the conditional Lindeberg condition holds for the martingale increments $\eta\xi_{t}(\theta_{t})$ (e.g., via a conditional CLT under exchangeability as in Section 4).
Remark 5.3 (Verifiable sufficient conditions).
Assumption 5.2 holds if $\sup_{\theta\in\mathcal{U}}\mathbb{E}[\|\psi(\theta;Z)\|_{2}^{2+\delta}]<\infty$ for some $\delta>0$ (guaranteed by Assumption 3.8(b)), $G^{\star}(\theta)$ is continuous on $\mathcal{U}$, and the mini-batch is drawn by simple random sampling from an exchangeable sequence.
Theorem 5.4 (Diffusion approximation).
Suppose Assumptions 5.15.2 hold. For fixed batch size $b\geq 1$ and constant step size $\eta>0$, define the effective temperature $\tau:=\eta/b$. For sufficiently small $\eta$, the SGD recursion admits a diffusion approximation: the piecewise-linear interpolation $\bar{\theta}^{\eta}(\cdot)$ converges weakly on $D([0,T];\mathbb{R}^{d})$ as $\eta\downarrow 0$ to the solution $\theta(\cdot)$ of the SDE
$d\theta_{s}=-h(\theta_{s})\,ds+\sqrt{\tau}\,C^{\star}(\theta_{s})\,dW_{s},\qquad C^{\star}(\theta)\,{C^{\star}(\theta)}^{\top}=G^{\star}(\theta),$ (27)
where $C^{\star}$ is, e.g., the symmetric PSD square root of $G^{\star}$,
where $W_{s}$ is standard $d$-dimensional Brownian motion. In particular, the diffusion has instantaneous covariance $\tau\,G^{\star}(\theta_{s})\,ds$. In the correctly specified likelihood case, $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$ at $\theta^{\star}$, so the diffusion covariance is locally Fisher-structured near the optimum.
@@
Step 1 (scaled interpolation). Write the recursion in signed-noise form $\theta_{t+1}=\theta_{t}-\eta h(\theta_{t})+\eta\,\zeta_{t+1}$ with $\zeta_{t+1}:=-\xi_{t}$, so $(\zeta_{t+1},\mathcal{F}_{t+1})$ is a martingale difference. Define the piecewise-linear interpolation $\bar{\theta}^{\eta}(s)$ on algorithmic time $s=t\eta$ by $\bar{\theta}^{\eta}(t\eta)=\theta_{t}$ with linear interpolation.
Step 2 (drift convergence). By local Lipschitzness of $h$ (Assumption 5.1) and the localization condition, $\sum_{k<s/\eta}\eta\,h(\theta_{k})\to\int_{0}^{s}h(\theta(u))\,du$ uniformly on compacts, by standard ODE-approximation theory (Kushner and Yin, 2003, Chapter 5).
Step 3 (predictable quadratic variation). Define $M^{\eta}(s):=\sum_{k<s/\eta}\eta\,\zeta_{k+1}$. By the alignment theorem (Theorem 4.3), $\mathbb{E}[\zeta_{k+1}\zeta_{k+1}^{\top}\mid\mathcal{F}_{k}]=\frac{1}{b}\,G^{\star}(\theta_{k})+R_{k}$ with $\|R_{k}\|_{\operatorname{op}}\leq(c/b)\rho_{k}$ and $\rho_{k}\to 0$ locally. Using $\eta^{2}/b=\eta\tau$:
@@
$\langle M^{\eta}\rangle(s)=\sum_{k<s/\eta}\eta^{2}\mathbb{E}[\zeta_{k+1}\zeta_{k+1}^{\top}\mid\mathcal{F}_{k}]=\tau\sum_{k<s/\eta}\eta\,G^{\star}(\theta_{k})+o(1)\;\to\;\tau\!\int_{0}^{s}\!G^{\star}(\theta(u))\,du$
in probability, uniformly on $[0,S]$.
Step 4 (Lindeberg and tightness). The fourth moment bound (Assumption 3.8(b)) yields $\sum_{k<S/\eta}\mathbb{E}[\|\eta\zeta_{k+1}\|^{2}\mathbf{1}\{\|\eta\zeta_{k+1}\|>\varepsilon\}\mid\mathcal{F}_{k}]\leq CS\eta^{3}m_{4}/(\varepsilon^{2}b^{2})\to 0$, verifying the conditional Lindeberg condition. Tightness of $\{\bar{\theta}^{\eta}\}$ in $D([0,S];\mathbb{R}^{d})$ follows from Aldous’ criterion and the Lipschitz-controlled drift.
@@
Step 5 (martingale FCLT). By the martingale central limit theorem (Ethier and Kurtz, 1986, Theorem 7.1.4), $M^{\eta}(\cdot)\Rightarrow M(\cdot)$ weakly in $D([0,S];\mathbb{R}^{d})$ where $M$ is a continuous martingale with quadratic variation $\tau\int_{0}^{s}G^{\star}(\theta(u))\,du$. Lévy’s characterization gives a Brownian motion $W$ such that $M(s)=\int_{0}^{s}\sqrt{\tau}\,C^{\star}(\theta(u))\,dW_{u}$. Combining drift and martingale convergence identifies the limit as the unique solution of (27). ∎
5.3 The OU regime near a nondegenerate critical point
Assume $\theta^{\star}$ is a stationary point: $h(\theta^{\star})=0$, and define the local curvature $H^{\star}:=\nabla h(\theta^{\star})=\nabla^{2}L(\theta^{\star})$. Let $\Delta_{s}:=\theta_{s}-\theta^{\star}$ denote the centered diffusion.

### 15 Apr 2026 v4

@@ 3.7 Key geometric parameters (preview)
4 Covariance Identification for Mini-Batch Noise
@@ 4.1 Finite-population covariance identity (exact, conditional on data)
4.2 From exchangeability to the population covariance
@@
Theorem 4.3 (Covariance identification under exchangeability).
@@
When the conditional mean $\mathbb{E}[\psi(\theta;Z)\mid\mu]$ does not vary with $\mu$ (identifiability under a single data-generating distribution), the unconditional covariance reads $\mathrm{Cov}(g_{B}(\theta))=b^{-1}G(\theta)$ where $G(\theta):=\mathbb{E}_{\mu}[G_{\mu}(\theta)]$; after projection onto the identifiable tangent space (Section 3), we denote this $G^{\star}(\theta)=\Pi_{\theta}G(\theta)\Pi_{\theta}^{\top}$. The remainder of the paper works with $G^{\star}$. Interpretation. Equation (21) is exact (not an approximation) when one marginalizes over both the mini-batch index $B$ and the dataset realization conditional on $\mu$. The finite-population correction (1b/n)(1-b/n) of Lemma 4.1 is present when one conditions on a fixed dataset, but it disappears at the population level because the randomness of the dataset itself contributes a complementary term through the total-covariance decomposition. Equation (22) below makes this bookkeeping explicit.
@@
Figure 1: Isotropic noise (left, dashed circle) vs. Fisher-aligned noise (right, solid ellipse) at the same total variance. Batch size rescales temperature but not shape.
@@
5 Diffusion Limits: Fluid Scale, Fluctuation Scale, and the OU Regime
This section links the discrete-time SGD recursion to a continuous-time stochastic process whose diffusion coefficient is determined by the mini-batch covariance $G^{\star}(\theta)$ identified in Section 4. The link is not a raw diffusion limit at fixed batch size: the martingale quadratic variation of the raw interpolation is of order $\eta/b$ and vanishes as $\eta\downarrow 0$, so the raw process converges to the deterministic gradient-flow ODE. The nondegenerate stochastic limit appears only after centering around that ODE and rescaling by $\sqrt{b/\eta}$, yielding a fluctuation process whose limiting covariance structure is exactly $G^{\star}$. Near a nondegenerate optimum, this fluctuation process is an Ornstein–Uhlenbeck (OU) process whose stationary covariance $\Sigma_{U}$ satisfies a fluctuation-scale Lyapunov equation. On the original iterate scale, the linearized discrete recursion of Proposition 5.13 admits a unique stationary covariance $\Sigma_{\eta}^{\mathrm{lin}}$ that matches $(\eta/b)\Sigma_{U}$ at leading order in $\eta$, recovering the paper’s “effective temperature” $\tau=\eta/b$ on the correct scale; we do not claim an invariant law for the nonlinear SGD iterates. Diffusion approximations are a standard tool in OR for reducing stochastic systems to tractable Brownian models (Harrison, 1985; Whitt, 2002); the present setting is analogous, with the sampling mechanism playing the role of the arrival process.
Throughout this section we work in the fresh-sampling / population conditioning regime of Theorem 4.3: at each iterate the mini-batch is a simple random sample drawn with fresh randomness from the population (equivalently, conditional on the de Finetti directing measure $\mu$). The fixed-dataset reuse case with (1b/n)(1-b/n) correction belongs to Lemma 4.1 and Proposition 4.5 and is not used below.
5.1 SGD recursion and the martingale-noise decomposition
Fix a batch size $b\in\mathbb{N}$ and a step size $\eta>0$. Let
@@ 5.1 SGD recursion and the martingale-noise decomposition
and write the mini-batch gradient estimator as $g_{t}(\theta_{t})=h(\theta_{t})+\xi_{t}(\theta_{t})$ with $\mathbb{E}[\xi_{t}(\theta_{t})\mid\mathcal{F}_{t}]=0$, where $\mathcal{F}_{t}$ is the natural filtration. The SGD recursion is
$\theta_{k+1}^{\eta}=\theta_{k}^{\eta}-\eta\,h(\theta_{k}^{\eta})-\eta\,\xi_{k+1}^{\eta},$ (23)
with the superscript $\eta$ emphasizing the dependence of the iterate path on the step size. Under the fresh-sampling conditioning,
$\mathbb{E}\!\bigl[\xi_{k+1}^{\eta}(\xi_{k+1}^{\eta})^{\top}\mid\mathcal{F}_{k}^{\eta}\bigr]\;=\;\frac{1}{b}\,G^{\star}(\theta_{k}^{\eta})+r_{k+1}^{\eta},$ (24)
with remainder $r_{k+1}^{\eta}$ satisfying the uniform bound stated in Assumption 5.2 below. We call $\tau:=\eta/b$ the effective temperature:
@@
the scalar that sets the size of the stationary fluctuations on the original iterate scale (Corollary 5.11).
The covariance identification of Section 4 determines the noise geometry, but at fixed batch size $b$ the raw interpolation of SGD on algorithmic time $t=k\eta$ has martingale quadratic variation of order $\eta/b$, which vanishes as $\eta\downarrow 0$. Accordingly, the raw process converges to the deterministic gradient-flow ODE. The correct nondegenerate stochastic limit is obtained by centering around that ODE and rescaling by $\sqrt{b/\eta}$, yielding a fluctuation process whose limiting covariance structure is exactly $G^{\star}$.
5.2 Assumptions for the fluid and fluctuation limits
Assumption 5.1 (Local regularity and ODE flow).
There is an open neighborhood $\mathcal{U}\subseteq\mathbb{R}^{d}$ such that:
@@ 5.2 Assumptions for the fluid and fluctuation limits
    $h\in C^{1}(\mathcal{U};\mathbb{R}^{d})$ and $\nabla h$ is locally Lipschitz on $\mathcal{U}$;
@@
    $G^{\star}:\mathcal{U}\to\mathbb{S}_{+}^{d}$ is continuous;
@@
    the ODE $\dot{\bar{\theta}}(t)=-h(\bar{\theta}(t))$ with $\bar{\theta}(0)=\theta_{0}$ has a unique solution on $[0,T]$ remaining in a compact set $K\Subset\mathcal{U}$.
Assumption 5.2 (Martingale noise and conditional covariance).
The sequence $(\xi_{k+1}^{\eta},\mathcal{F}_{k+1}^{\eta})$ is a martingale difference, and the conditional covariance admits the expansion (24) with
$\sup_{k\leq T/\eta}\|r_{k+1}^{\eta}\|_{\operatorname{op}}\;\xrightarrow{\,\mathbb{P}\,}\;0\qquad\text{as }\eta\downarrow 0.$
Assumption 5.3 (Conditional fourth-moment bound on $K$).
There is a constant $C_{K}<\infty$ such that
$\sup_{\eta>0}\,\sup_{k\leq T/\eta}\,\mathbb{E}\!\left[\|\xi_{k+1}^{\eta}\|^{4}\,\mathbf{1}_{\{\theta_{k}^{\eta}\in K\}}\,\big|\,\mathcal{F}_{k}^{\eta}\right]\;\leq\;C_{K}\quad\text{a.s.}$
Assumption 5.4 (Localization).
$\Pr\!\bigl(\theta_{k}^{\eta}\in K\text{ for all }k\leq T/\eta\bigr)\to 1$ as $\eta\downarrow 0$.
Assumption 5.5 (Initial fluctuations converge).
$\theta_{0}^{\eta}\to\theta_{0}$, and $U_{0}^{\eta}:=\sqrt{b/\eta}\,(\theta_{0}^{\eta}-\theta_{0})\Rightarrow U_{0}$ for some random variable $U_{0}$.
Remark 5.6 (Verifiability).
Assumptions 5.25.3 hold if $\sup_{\theta\in\mathcal{U}}\mathbb{E}[\|\psi(\theta;Z)\|^{4}]<\infty$ (guaranteed by Assumption 3.8(b)), $G^{\star}$ is continuous on $\mathcal{U}$, and the mini-batch is drawn by simple random sampling from an exchangeable sequence. Assumption 5.4 is standard in stochastic approximation (Kushner–Yin style) and holds, e.g., when the ODE is globally stable and the initial iterate is in the basin.
5.3 Fluid limit and fluctuation theorem
Theorem 5.7 (Fluid limit and fluctuation limit at fixed batch size).
Fix $T>0$ and $b\in\mathbb{N}$, and suppose Assumptions 5.15.5 hold. Let $\hat{\theta}^{\eta}(t):=\theta_{\lfloor t/\eta\rfloor}^{\eta}$ be the step interpolation of the SGD iterates (23), and let $\bar{\theta}$ solve the ODE in Assumption 5.1(iii). Then:
  1. (a)
    Fluid limit.
    $\sup_{t\in[0,T]}\bigl\|\hat{\theta}^{\eta}(t)-\bar{\theta}(t)\bigr\|\;\xrightarrow{\,\mathbb{P}\,}\;0.$
  2. (b)
    Fluctuation FCLT. The rescaled fluctuation process
    $U^{\eta}(t)\;:=\;\sqrt{\tfrac{b}{\eta}}\,\bigl(\hat{\theta}^{\eta}(t)-\bar{\theta}(t)\bigr)$
    converges weakly in $D([0,T];\mathbb{R}^{d})$ (equivalently, in $C([0,T];\mathbb{R}^{d})$, as the limit is continuous) to the unique weak solution $U$ of the linear SDE
[…]
    where $W$ is a standard $d$-dimensional Brownian motion and $C^{\star}(\theta){C^{\star}(\theta)}^{\top}=G^{\star}(\theta)$ (e.g., the symmetric PSD square root of $G^{\star}$).
Equivalently, on the original iterate scale,
$\hat{\theta}^{\eta}(t)\;=\;\bar{\theta}(t)\;+\;\sqrt{\tfrac{\eta}{b}}\,U_{t}\;+\;o_{\mathbb{P}}\!\Bigl(\sqrt{\tfrac{\eta}{b}}\Bigr),$ (27)
uniformly on $[0,T]$.
@@ 5.3 Fluid limit and fluctuation theorem
We follow the standard stochastic-approximation FCLT architecture (cf. Kushner and Yin, 2003, Ch. 8), tailored to the fresh-sampling covariance (24). Write $t_{k}:=k\eta$ and $\bar{\theta}_{k}:=\bar{\theta}(t_{k})$; on the localization event of Assumption 5.4, which has probability tending to $1$, the iterates stay in $K$ for all $k\leq T/\eta$, and we work under this event without further comment.
Step 1 (fluid limit). Define the raw error $e_{k}^{\eta}:=\theta_{k}^{\eta}-\bar{\theta}_{k}$ and the ODE local truncation error $\rho_{k+1}^{\eta}:=\bar{\theta}_{k+1}-\bar{\theta}_{k}+\eta\,h(\bar{\theta}_{k})$. Local Lipschitzness of $h$ on $K$ gives $\sup_{k\leq T/\eta}\|\rho_{k+1}^{\eta}\|\leq C_{T}\eta^{2}$ for some $C_{T}<\infty$. Subtracting the Euler step for $\bar{\theta}$ from (23),
$e_{k+1}^{\eta}\;=\;e_{k}^{\eta}\;-\;\eta\bigl(h(\theta_{k}^{\eta})-h(\bar{\theta}_{k})\bigr)\;-\;\eta\,\xi_{k+1}^{\eta}\;-\;\rho_{k+1}^{\eta}.$ (28)
Let $N_{n}^{\eta}:=\eta\sum_{j=0}^{n-1}\xi_{j+1}^{\eta}$; this is an $\mathcal{F}_{k}^{\eta}$-martingale. Set $\bar{G}_{K}:=\sup_{\theta\in K}\mathrm{Tr}G^{\star}(\theta)<\infty$ (finite by continuity of $G^{\star}$ on the compact set $K$). By Doob’s maximal inequality and Assumptions 5.25.3,
$\mathbb{E}\!\Bigl[\sup_{n\leq T/\eta}\|N_{n}^{\eta}\|^{2}\Bigr]\;\leq\;4\,\eta^{2}\sum_{j<T/\eta}\mathbb{E}\|\xi_{j+1}^{\eta}\|^{2}\;\leq\;4\,\eta^{2}\cdot\frac{T}{\eta}\cdot\frac{\bar{G}_{K}+o(1)}{b}\;=\;O(\eta/b).$
Hence $\sup_{n\leq T/\eta}\|N_{n}^{\eta}\|\xrightarrow{\mathbb{P}}0$. Using local Lipschitzness of $h$ on $K$ (constant $L_{K}$) in (28),
$\|e_{n}^{\eta}\|\;\leq\;\|e_{0}^{\eta}\|\;+\;L_{K}\eta\sum_{j=0}^{n-1}\|e_{j}^{\eta}\|\;+\;\sup_{m\leq n}\|N_{m}^{\eta}\|\;+\;C_{T}\,T\,\eta,$
and discrete Grönwall gives
$\sup_{n\leq T/\eta}\|e_{n}^{\eta}\|\;\leq\;e^{L_{K}T}\Bigl(\|e_{0}^{\eta}\|+\sup_{m\leq T/\eta}\|N_{m}^{\eta}\|+C_{T}T\eta\Bigr)\;\xrightarrow{\,\mathbb{P}\,}\;0,$
since $\|e_{0}^{\eta}\|\to 0$ by Assumption 5.5. This establishes part (a).
Step 2 (rescaled recursion). Set $U_{k}^{\eta}:=\sqrt{b/\eta}\,e_{k}^{\eta}$. Local Lipschitzness of $\nabla h$ on $K$ gives the Taylor expansion $h(\theta_{k}^{\eta})-h(\bar{\theta}_{k})=A_{k}^{\eta}e_{k}^{\eta}+q_{k}^{\eta}$ with $A_{k}^{\eta}:=\nabla h(\bar{\theta}_{k})$ and $\|q_{k}^{\eta}\|\leq L^{\prime}_{K}\|e_{k}^{\eta}\|^{2}$. Multiplying (28) by $\sqrt{b/\eta}$ and using $\eta\sqrt{b/\eta}=\sqrt{\eta b}$,
$U_{k+1}^{\eta}\;=\;U_{k}^{\eta}\;-\;\eta\,A_{k}^{\eta}\,U_{k}^{\eta}\;+\;\Delta M_{k+1}^{\eta}\;+\;\varepsilon_{k+1}^{\eta},$ (29)
with
$\Delta M_{k+1}^{\eta}\;:=\;-\sqrt{\eta b}\,\xi_{k+1}^{\eta},\qquad\varepsilon_{k+1}^{\eta}\;:=\;-\sqrt{\eta b}\,q_{k}^{\eta}\;-\;\sqrt{\tfrac{b}{\eta}}\,\rho_{k+1}^{\eta}.$ (30)
Step 3 (tightness scale and remainders). From Step 1, $\sup_{n}\|e_{n}^{\eta}\|=e^{L_{K}T}(\|e_{0}^{\eta}\|+\sup\|N^{\eta}\|+C_{T}T\eta)$. Multiplying by $\sqrt{b/\eta}$,
$\sup_{n}\|U_{n}^{\eta}\|\;\leq\;e^{L_{K}T}\Bigl(\|U_{0}^{\eta}\|+\sqrt{\tfrac{b}{\eta}}\sup_{m}\|N_{m}^{\eta}\|+C_{T}T\sqrt{b\eta}\Bigr).$
The first term is tight by Assumption 5.5; $\mathbb{E}[(b/\eta)\sup_{m}\|N_{m}^{\eta}\|^{2}]\leq 4b/\eta\cdot O(\eta/b)=O(1)$, so the second is $O_{\mathbb{P}}(1)$; the third tends to 0. Hence
$\sup_{n\leq T/\eta}\|U_{n}^{\eta}\|=O_{\mathbb{P}}(1).$ (31)
Using (31) in the Taylor-remainder bound,
$\Bigl\|\sum_{k<n}\sqrt{\eta b}\,q_{k}^{\eta}\Bigr\|\;\leq\;\sqrt{\eta b}\cdot L^{\prime}_{K}\sum_{k<n}\|e_{k}^{\eta}\|^{2}\;\leq\;L^{\prime}_{K}\,T\,\sqrt{\tfrac{\eta}{b}}\,\sup_{k}\|U_{k}^{\eta}\|^{2}\xrightarrow{\mathbb{P}}0,$
and, since $\|\rho_{k+1}^{\eta}\|\leq C_{T}\eta^{2}$,
$\Bigl\|\sum_{k<n}\sqrt{\tfrac{b}{\eta}}\,\rho_{k+1}^{\eta}\Bigr\|\;\leq\;\frac{T}{\eta}\,C_{T}\eta^{2}\sqrt{\tfrac{b}{\eta}}\;=\;C_{T}T\sqrt{b\eta}\to 0.$
Hence $\sup_{t\leq T}\bigl\|\sum_{k<t/\eta}\varepsilon_{k+1}^{\eta}\bigr\|\xrightarrow{\mathbb{P}}0$.
Step 4 (martingale FCLT). Let $M^{\eta}(t):=\sum_{k<t/\eta}\Delta M_{k+1}^{\eta}=-\sqrt{\eta b}\sum_{k<t/\eta}\xi_{k+1}^{\eta}$. By (24) its predictable quadratic variation is
@@
$\langle M^{\eta}\rangle(t)\;=\;\eta b\sum_{k<t/\eta}\!\mathbb{E}\!\bigl[\xi_{k+1}^{\eta}(\xi_{k+1}^{\eta})^{\top}\mid\mathcal{F}_{k}^{\eta}\bigr]\;=\;\eta\sum_{k<t/\eta}G^{\star}(\theta_{k}^{\eta})\;+\;\eta b\sum_{k<t/\eta}r_{k+1}^{\eta}.$
The remainder is bounded uniformly in $t\leq T$ by $Tb\cdot\sup_{k}\|r_{k+1}^{\eta}\|\xrightarrow{\mathbb{P}}0$ under Assumption 5.2. For the leading term, continuity of $G^{\star}$ on $K$ and the fluid limit $\sup_{k}\|\theta_{k}^{\eta}-\bar{\theta}_{k}\|\to 0$ (Step 1) give $\eta\sum_{k<t/\eta}\|G^{\star}(\theta_{k}^{\eta})-G^{\star}(\bar{\theta}_{k})\|\to 0$ in probability uniformly in $t$; and $\eta\sum_{k<t/\eta}G^{\star}(\bar{\theta}_{k})\to\int_{0}^{t}G^{\star}(\bar{\theta}(s))\,ds$ is a Riemann-sum approximation with continuous integrand. Hence
$\langle M^{\eta}\rangle(t)\;\Rightarrow\;\int_{0}^{t}G^{\star}(\bar{\theta}(s))\,ds\qquad\text{uniformly on }[0,T].$
For the conditional Lindeberg condition, using $\mathbf{1}_{\{x>\varepsilon\}}\leq x^{2}/\varepsilon^{2}$,
$\sum_{k<T/\eta}\!\mathbb{E}\!\bigl[\|\Delta M_{k+1}^{\eta}\|^{2}\mathbf{1}_{\{\|\Delta M_{k+1}^{\eta}\|>\varepsilon\}}\mid\mathcal{F}_{k}^{\eta}\bigr]\;\leq\;\frac{(\eta b)^{2}}{\varepsilon^{2}}\sum_{k<T/\eta}\mathbb{E}[\|\xi_{k+1}^{\eta}\|^{4}\mid\mathcal{F}_{k}^{\eta}]\;\leq\;\frac{\eta\,b^{2}\,T\,C_{K}}{\varepsilon^{2}}\to 0.$
@@
The martingale FCLT (Ethier and Kurtz, 1986, Theorem 7.1.4) then gives $M^{\eta}\Rightarrow M$ in $D([0,T];\mathbb{R}^{d})$, where $M$ is a continuous Gaussian martingale with quadratic variation $\int_{0}^{\cdot}G^{\star}(\bar{\theta}(s))\,ds$. Lévy’s characterization yields a Brownian motion $W$ with $M_{t}=\int_{0}^{t}C^{\star}(\bar{\theta}(s))\,dW_{s}$, and $C^{\star}(\bar{\theta}(\cdot))$ can be taken as the (continuous) symmetric PSD square root of $G^{\star}(\bar{\theta}(\cdot))$.
Step 5 (identify the limiting SDE). Summing (29) and using Step 3,
$U^{\eta}(t)\;=\;U_{0}^{\eta}\;-\;\int_{0}^{t}\!\nabla h\!\bigl(\bar{\theta}(\lfloor s/\eta\rfloor\eta)\bigr)\,U^{\eta}(s)\,ds\;+\;M^{\eta}(t)\;+\;o_{\mathbb{P}}(1),$
uniformly on $[0,T]$. Continuity of $s\mapsto\nabla h(\bar{\theta}(s))$ and tightness $\sup_{t}\|U^{\eta}(t)\|=O_{\mathbb{P}}(1)$ give $\int_{0}^{t}\nabla h(\bar{\theta}(\lfloor s/\eta\rfloor\eta))U^{\eta}(s)ds-\int_{0}^{t}\nabla h(\bar{\theta}(s))U^{\eta}(s)ds\to 0$ uniformly in probability. Tightness of $\{U^{\eta}\}$ in $D([0,T];\mathbb{R}^{d})$ follows from (31) and the controlled drift together with Aldous’ criterion applied to the martingale term (whose FCLT is in hand). Any subsequential weak limit $U$ therefore satisfies the stochastic integral equation
$U_{t}\;=\;U_{0}\;-\;\int_{0}^{t}\nabla h(\bar{\theta}(s))\,U_{s}\,ds\;+\;\int_{0}^{t}C^{\star}(\bar{\theta}(s))\,dW_{s},$
which is (26) in integrated form. The linear SDE (26) with continuous coefficients admits a unique weak solution, so the entire sequence converges: $U^{\eta}\Rightarrow U$.
Step 6 (expansion). By construction $\hat{\theta}^{\eta}(t)-\bar{\theta}(t)=\sqrt{\eta/b}\,U^{\eta}(t)$; combining (a) and (b) gives (27). The asymptotic equivalence of step and piecewise-linear interpolations under Assumption 5.3 is standard and yields the same limit for either choice. ∎
Remark 5.8 (Why the raw process is deterministic in the limit).
At fixed $b$, the martingale variation of the raw interpolation is $\sum_{k<t/\eta}\eta^{2}\,b^{-1}G^{\star}(\theta_{k}^{\eta})\sim(\eta/b)\int_{0}^{t}G^{\star}(\bar{\theta}(s))\,ds$, which vanishes as $\eta\downarrow 0$. Hence the raw stochastic fluctuation is invisible at the original scale, and a nondegenerate Gaussian limit is recovered only after centering and rescaling by $\sqrt{b/\eta}$. This is the standard stochastic-approximation functional CLT regime; the diffusion lives on the fluctuation scale, not in the raw iterate path.
5.4 The OU regime as the equilibrium specialization
Suppose $\theta^{\star}$ is a stationary point: $h(\theta^{\star})=0$, and $\theta_{0}=\theta^{\star}$ so that the ODE flow is trivial, $\bar{\theta}(t)\equiv\theta^{\star}$. Then the drift coefficient in (26) is the constant matrix $H^{\star}:=\nabla h(\theta^{\star})=\nabla^{2}L(\theta^{\star})$, and the diffusion coefficient is the constant matrix $C^{\star}(\theta^{\star})$.